Shift Summary, EVE Nov 15, 1999

Ricardo_Barbosa (rbarbosa@fnal.gov)
Tue, 16 Nov 1999 01:13:26 -0600

Shift Summary for EVE Shift, 15/Nov/99

On Shift: Ricardo, RickT

Hours of beam available at all intensities: 8.4 hrs
Hours of beam available at nominal intensity: 8.0 hrs
Hours of E799 data (special runs not included): 6.0 hrs
Hours of special runs (laser scans, muon runs...): 0.1 hrs

Accounting of lost time:

tape changes. initializing runs...

========================================================================
Runs ended this shift:
Run 14982 Beam_799 tapes full
Run 14983 muon_run


Tapes brought to FCC this shift: UPA-C 142-144, UPS 048

========================================================================
Accesses during shift: one
Investigating low voltage sector-9 failure

========================================================================
Overview of shift:

The problem with this shift: a few minor problems (see below).
Beam intensity low (4E12) about half hour.

========================================================================
Description of Shift:

run spill time comment

------------------------------------------------------------------------

15:30 Julie + Sydney on access to replace CsI
LV supply. BobT added to controlled access
some time later.

17:00 Cosmic run after power supply swap

14982 17:09 Start a new run!

18:15 Some wide SOD's found

19:15 Problem with RC. (see logbook)

19:40 Strange behavior for a cosmic telescope...
low illumination (see logbook)

20:30 Problems with Etot (threshold 4)

21:12 Running at 8.5E12. MCR called to tweak up
intensity

21:26 Kumquat problems (see logbook)

21:40 Regenerator counting when we don't expect.
RickT called Elizabeth and REG count was turn off.

209-210 19:40 Mainframe 15 turned off. Accidental counters
missing
for half of these spills.

250-251 22:50 NM2SEED2 alarm. But all right at the monitor...

23:02 Beam intensity decreasing over the last half
hour...

23:19 Phoned MCR.

23:21 Intensity back up.

14983 23:44 Trigger in error state. Reload kumquat,
run kfrend and go smoothly...

14984 0:10 Start run!!!

========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift